OpenStudy (anonymous):

3 ( d + 12 ) = 8 - 4d

OpenStudy (anonymous):

3d+36=8-4d 3d+4d=36-8 7d=28 therefored=28/7=4

hartnn (hartnn):

3d+4d = 8-36 7d=-28 d=-28/7=-4